OpenStudy (anonymous):

Evaluate the following expression: cos((cos^-1(4)/7)-((sin^-1(2))/12)) Give an exact answer

OpenStudy (anonymous):

No trig functions allowed.

OpenStudy (idku):

Well, \(\cos^{-1}\left\{\cos(\theta)\right\}=\theta \)

OpenStudy (idku):

I meant \(\cos\left\{\cos^{-1}(\theta)\right\}=\theta\)

OpenStudy (idku):

\(\cos\left\{\dfrac{\cos^{-1}(4)}{7}-\dfrac{\cos^{-1}(2)}{12}\right\} \) this?

OpenStudy (idku):

oh the second one is sin^-1

OpenStudy (idku):

\(\cos\left\{\dfrac{\cos^{-1}(4)}{7}-\dfrac{\sin^{-1}(2)}{12}\right\} \)

OpenStudy (anonymous):

Yeah

OpenStudy (idku):

I will do an example

OpenStudy (anonymous):

ok

OpenStudy (idku):

I had to leave, apolgoize

OpenStudy (anonymous):

No Problem